1. In a cricket match, a batswoman hits a boundary 6 times out of 30 balls she plays. Find the probability that she did not hit a boundary.

Grade:12th pass

2 Answers

Pawan Prajapati
3 years ago
lbtal number of balls = 30. NWQber of hitting the boundary = 6. Number of balls when boundary is not bit = 30 ..,. 6 = 24. Probability that she did not bit the boundary = = .
Debarati
2 years ago
No. of times she didn’t hit the boundary= 30-6= 24
Probability of not hitting the boundary= 24/30= 12/15

I hope you understood.